Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-023_2000-029

Mission Operations Days: 2000/023 to 2000/029
January 23, 2000 00:00:00 GMT to January 29, 2000 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
Terra (AM-1) has an anomaly with the Propulsion Thrusters (that is still being worked)
MODIS is in Safe Mode
MODIS has no known Anomalies
ComponentState(s)Comment
BlackbodyA Off; B OffNominal
Calibration ElectronicsA Off; B OffNominal
Control ProcessorA Off; B OnNominal
Nadir DoorUnlatched, ClosedNominal
Space View DoorUnlatched, ClosedNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA Off; B OffNominal
FDDI PortA Off; B OffNominal
FIFO Memory#1 & #2 Off; #3 & #4 OffNominal
Format ProcessorA Off; B OffNominal
Power Supplies#1 Off, #2 OnNominal
PV VIS FPAA Off; B OffNominal
PV NIR FPAA Off; B OffNominal
PV SMIR FPAA Off; B OffNominal
PV LWIR FPAA Off; B OffNominal
PC LWIR FPAA Off; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOffNominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA Off; B OffNominal
SDSMOffNominal
SRCAOffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BDNominal
Inhibit Ids Set31,32,33,53,58BB, CE, CP, SDSM, SD
TMONs EnabledNoneNominal

This Week’s Completed MODIS Activities:

Sunday, January 23, 2000

2000/023 21:32 Turned on CS and IS OG heaters
2000/023 21:35 Turned on OS OG heater
2000/023 21:37 Opened the Space View Door

Monday, January 24, 2000

2000/024 14:18 Turned off the outgas heaters
2000/024 18:58 Transitioned to Science Mode on the B side
2000/024 19:40 Performed PV electronics calibration
2000/024 22:05 Turn on blackbody to 285K to warm for A side electronics

Tuesday, January 25, 2000

2000/025 16:05 Memory Load test
2000/025 17:10 Memory Dump test
2000/025 15:45 MODIS FPAs started controllling at 83K
2000/025 18:26 Commanded to Safe Mode by Spacecraft during burn

Wednesday, January 26, 2000

None

Thursday, January 27, 2000

None

Friday, January 28, 2000

None

Saturday, January 29, 2000

None

This Week's Scheduled MODIS Activities Not Completed:

The following activities were scheduled to occur, however were postponed due to the Solid State Recorder failure and the transition to Safe Mode.

2000/025 Jan 25th, 2000 Turn off PV and PC DC Restore (~12:09 EST)
2000/025 Jan 25th, 2000 Set formatter encoder delta to –3072 (~12:09 EST)
2000/025 Jan 25th, 2000 View SV with EV during two moon in SV events
2000/025 Jan 25th, 2000 Set formatter encoder delta to 0 (~15:13 EST)
2000/025 Jan 25th, 2000 Turn on PV and PC DC Restore (~15:13 EST)
2000/025 Jan 25th, 2000 Open SDD to Screened and then close SDD (~16:43 EST)
2000/025 Jan 25th, 2000 Open SDD to Open and then close SDD (~18:40 EST)

2000/026 Jan 26th, 2000 Turn off MODIS
2000/026 Jan 26th, 2000 Turn on MODIS on the A side
2000/026 Jan 26th, 2000 Transittion to Science Mode on the A side
2000/026 Jan 26th, 2000 Set Blackbody duty cycle to FULL
2000/026 Jan 26th, 2000 Perform PV and PC electronics calibrations (ATC)
2000/026 Jan 26th, 2000 Perform SDSM calibrations w/ SD Open and Screened Utilizes expedited data.
2000/026 Jan 26th, 2000 Perform BB calibration to 315K (ATC)

Upcoming MODIS Events:

MODIS will recover from Safe Mode once the MODIS team is comfortable that the instrument won’t be back in Safe Mode in the near future. The schedule for this transition will be discussed once several orbit attainment burns have been successfully completed.

MODIS Anomalies:

Many transient red and yellow alarms were triggered during the transition to science mode. The limit violation which took the longest to clear was for MOD_TA_PV_SM_PWB5_11. This temperature sensor is noticably colder (~4 to 12K) than the other 5 SAM MUX temperatures.

Also during the transition to Science Mode, the command and data handling people received two errors related to MODIS; SFE data overflow errors and taxi errors. In addition, EDOS noticed 89 time jumps in the science data. Both the SFE and time jump errors lasted until approximately 10 minutes after the science mode transition procedure had completed.

General Instrument Comments:

MODIS is currently in Safe Mode with all three doors closed and unlatched. The timing generator is powered on the B side due to the transition to Safe Mode from Science Mode.

MODIS “science data” was transmitted to ground based antennas via the X-band transponder. This science data is not “real” because the MODIS formatter is not powered on and the nadir aperture door is closed.

The load and dump of a sample macro 31 table were successful. In addition, a dump of table 18 (most recent door position trigger values) was dumped to verify the number of steps the space view door moved during the open and close processes.

MODIS Telemetry Trends:

MODIS focal planes started controlling at 83K approximately 45 hours after the transition to Science Mode.

Non-MODIS Significant Events:

2000/025: An SSR failure prevented further capture of MODIS science data. No further data was collected prior to transition to Safe Mode. Approximately 6 hours of the 9 hours of science data produced before the SSR anomaly occurred were captured and distributed to the GDAAC.

2000/026: Spacecraft and instruments sent to Safe Mode due to anomaly during 1.024 second burn.

Analysis and discussion of the thruster data continues. The burn schedule continues to be in a state of flux.

Limited Life Item Status:

The Space View Door was closed. All limited life items are well within lifetime ranges, although precise statistics for each item are still pending.